OpenStudy (anonymous):

How is sin(pi/2 + x) equal to cos x ??

terenzreignz (terenzreignz):

Magic, of course :)

OpenStudy (anonymous):

Type it in to your calc

OpenStudy (nathan917):

Yes. lol.

terenzreignz (terenzreignz):

Or, the fact that \[\Large \sin(\alpha+\beta) = \sin(\alpha)\cos(\beta)+\cos(\alpha)\sin(\beta)\] Pick whichever reason you like :D

OpenStudy (nathan917):

Or http://www.youtube.com/watch?v=n5RpLuWnCTs.

OpenStudy (anonymous):

there is no button of x in the calculator @Ninagf99

OpenStudy (anonymous):

in t 84 silver edition plus there is

terenzreignz (terenzreignz):

Identities, @koli123able you should study them well :) Plug in \[\alpha = \frac\pi 2\] and \[\beta = x\] And work out the magic :D